# Small Language Models in the Real World: Insights from Industrial Text Classification

### v1

@@ Abstract
With the emergence of ChatGPT, Transformer models have significantly advanced text classification and related tasks. Decoder-only models such as Llama exhibit strong performance and flexibility, yet they suffer from inefficiency on inference due to token-by-token generation, and their effectiveness in text classification tasks heavily depends on prompt quality. Moreover, their substantial GPU resource requirements often limit widespread adoption. Thus, the question of whether smaller language models are capable of effectively handling text classification tasks emerges as a topic of significant interest. However, the selection of appropriate models and methodologies remains largely underexplored. In this paper, we conduct a comprehensive evaluation of prompt engineering and supervised fine-tuning methods for transformer-based text classification. Specifically, we focus on practical industrial scenarios, including email classification, legal document categorization, and the classification of extremely long academic texts. We examine the strengths and limitations of smaller models, with particular attention to both their performance and their efficiency in Video Random-Access Memory (VRAM) utilization, thereby providing valuable insights for the local deployment and application of compact models in industrial settings.  [^1] https://github.com/DobricLilujun/agentCLS/
@@ 2.2 Background
The earliest systematic studies included probabilistic model-based methods such as Naive Bayes Joachims (1998). He was the first to apply Support Vector Machines (SVM) to text classification tasks. With the advent of neural networks, early research primarily utilized embeddings and simple neural network architectures for text classification. Subsequently, Kim (2014) proposed a convolutional neural network-based approach for text classification, significantly improving classification performance at sentence-level feature extraction. In addition, generative classification models based on recurring neural networks (RNNs) have also shown remarkable performance, demonstrating greater robustness under distribution shifts Yogatama et al. (2017). However, they still struggle to effectively handle complex classification tasks that involve long texts Du et al. (2020). Later, the emergence of attention architectures led to extensive experimentation in various applications.
The advent of transformer-based architectures in 2018, notably BERT, brought about a paradigm shift in natural language classification tasks, yielding considerable performance enhancements Kora and Mohammed (2023); Pawar et al. (2024). Some knowledge distillation approaches Nityasya et al. (2022) have been explored to compress large BERT models into smaller, faster, and more efficient versions that can retain up to 97% of the original model’s classification performance. This observation has motivated our interest in directly using small open source models, which often achieve performance comparable to that of large models after distillation Zhu et al. (2024). For long-text classification, specialized bidirectional models such as Longformer Beltagy et al. (2020) and LegalBERT Chalkidis et al. (2020) have emerged in recent years, capable of handling ultralong documents and demonstrating excellent performance.
@@ 4.1 Main Performance
Additional models were used to validate the test set. However, ChatGPT was not evaluated on the IE dataset due to potential data leakage concerns. In contrast, Llama-3.3-70B-Instruct was run locally, allowing for GPU resource estimation and comprehensive metric evaluation. As presented in Table 3, the highest prompt engineering performance was achieved by ChatGPT-o1 mini. Meanwhile, in the IE dataset, which serves as our industrial database, an accuracy score of 0.800 was achieved by Llama-3.3-70B-Instruct. Regarding SLMs, we found the results particularly intriguing, especially in the context of prompt engineering. Given the relatively small size of these models, we did not expect them to achieve high performance. The final results for the 1B and 3B models aligned with our expectations, performing roughly at the level of random guessing. Interestingly, both the 3B and even the 1B models demonstrated a strong preference for few-shot prompting. This approach led to an improvement of over 10% compared to the base prompt on the EUR and LDD datasets, highlighting the importance of few-shot learning in the application of SLMs, as also emphasized in Brown et al. (2020a). Furthermore, we observed that both COD and COT provided limited improvements. In fact, on the LDD dataset, COD performed worse than COT and was nearly on par with the base prompt. Therefore, the use of COD and COT is not recommended as a solution for classification tasks in SLMs.
In the context of FT, we observed that SPT outperformed Prefix Tuning by a significant margin, although it also required substantially more training time. Prefix-tuning introduces a trainable part at every layer within the model, whereas SPT only incorporates a soft prompt at the input level. It is possible that SPT better preserves the model’s original language understanding, as it does not alter the overall architecture. In contrast, prefix-tuning’s modifications to the attention structure may disrupt the model’s inherent linguistic comprehension. Additionally, Fine-Tuning, which freezes all model parameters, exhibited the highest overall performance. Notably, ModernBERT achieved a performance of approximately 0.810 on the LDD dataset while requiring less training time and GPU memory, making it a promising candidate for academic English text classification. Limited exposure to French, other multilingual languages, and domain-specific corpora during training Warner et al. (2024) led to weaker performance on the IE dataset (primarily in French) and EUR (a domain-specific corpus).
@@ 4.2.3 Deeper Header?
In our primary experimental setting, we adhere to the definition of a “Header” as implemented in the Transformers library, referring to a single linear layer serving as the classification head. To further explore potential improvements using different levels of header, we experimented with replacing the standard single-layer header with a multi-layer linear architecture incorporating ReLU activations. Specifically, we constructed classification heads with 2 to 5 linear layers (hidden dimension = 256) and fine-tuned Llama-3.2-1B-Instruct model accordingly. As shown in Table 5, the results indicate that increasing the depth of the classification head yields only marginal gains, with performance plateauing beyond three layers. These findings suggest that deeper header architectures offer limited benefit in enhancing the classification accuracy or F1 score in this context,
@@ 4.4 Research Questions
For RQ1, “Can SLMs perform classification without any task-specific training?”, we found that text classification using SLMs faces several key challenges. Smaller models tend to exhibit limited logical reasoning capabilities and are more susceptible to generating hallucinations. Moreover, the performance ceiling is strongly influenced by the amount of available training data, while the intrinsic properties of the SLMs themselves also play a critical role in shaping classification outcomes.
@@
For RQ3, “How can the trade-off between computational efficiency and classification performance be optimized, and how can SLMs be more effectively deployed in practice?”, we found that although training the model consumes significant GPU resources, the SLMs are essentially unusable in their current form due to the lack of inference capability. We also tested Llama-3.3-70B-Instruct, which, although capable of achieving 80% accuracy in IE, still produces uncertain output. Therefore, FT transformers remains the only viable solution on SLMs which is portable and light weight. Finally, the limited capacity of SLMs creates a bottleneck on performance and the amount of labeled data also remains a key limitation. For real application, it is crucial to focus not only on data quality but also on the model’s inherent characteristics, such as multilingual comprehension. If resources are relatively abundant, opting for decoder-only models such as the Llama series would be a better choice.
@@ 5 Conclusion
In this study, we present a comprehensive evaluation of transformer-based models with a primary focus on text classification. We systematically investigate nearly all major approaches, including prompt engineering and supervised fine-tuning, with particular emphasis on lightweight models. A rigorous and standardized assessment is performed on widely adopted open-source models such as BERT and Llama (1B and 3B variants). Our experimental setup spans three benchmark datasets, including a real-world industrial scenario involving email history classification.

### v2

@@ Abstract
With the emergence of ChatGPT, Transformer models have significantly advanced text classification and related tasks. Decoder-only models such as Llama exhibit strong performance and flexibility, yet they suffer from inefficiency on inference due to token-by-token generation, and their effectiveness in text classification tasks heavily depends on prompt quality. Moreover, their substantial GPU resource requirements often limit widespread adoption. Thus, the question of whether smaller language models are capable of effectively handling text classification tasks emerges as a topic of significant interest. However, the selection of appropriate models and methodologies remains largely underexplored. In this paper, we conduct a comprehensive evaluation of prompt engineering and supervised fine-tuning methods for transformer-based text classification. Specifically, we focus on practical industrial scenarios, including email classification, legal document categorization, and the classification of extremely long academic texts. We examine the strengths and limitations of smaller models, with particular attention to both their performance and their efficiency in Video Random-Access Memory (VRAM) utilization, thereby providing valuable insights for the local deployment and application of compact models in industrial settings [^1] https://github.com/DobricLilujun/agentCLS/.
@@ 2.2 Background
The earliest systematic studies on text classification included probabilistic model-based methods such as Naive Bayes Joachims (1998). He was the first to apply Support Vector Machines (SVM) to text classification tasks. With the advent of neural networks, early research primarily utilized embeddings and simple neural network architectures for text classification. Subsequently, Kim (2014) proposed a convolutional neural network-based approach for text classification, significantly improving classification performance at sentence-level feature extraction. In addition, classification models based on Recurrent Neural Networks (RNNs) have also shown remarkable performance, demonstrating greater robustness under distribution shifts Yogatama et al. (2017). However, they still struggle to effectively handle complex scenarios in classification tasks such as long textsDu et al. (2020). Later, the emergence of attention architectures led to extensive experimentation in various applications.
The advent of transformer-based architectures in 2018, particularly BERT, brought about a paradigm shift in natural language classification tasks, resulting in considerable performance enhancements Kora and Mohammed (2023); Pawar et al. (2024). Some knowledge distillation approaches Nityasya et al. (2022) have also been explored to compress large BERT models into smaller, faster, and more efficient versions that can retain up to 97% of the original model’s classification performance. This observation has motivated our interest in directly using small open source models, which often achieve performance comparable to that of large models after distillation Zhu et al. (2024). For long text classification, specialized bidirectional models such as Longformer Beltagy et al. (2020) and LegalBERT Chalkidis et al. (2020) have emerged in recent years, capable of handling ultra-long documents and showing excellent performance. Nevertheless, their adoption in industry remains limited, primarily due to substantial GPU resource requirements and the need for custom CUDA kernels to support sliding-window attention, which also introduces compatibility challenges with the Huggingface Transformers framework.
@@ 4.1 Main Performance
Additional models were used to validate the test set in order to provide a reference performance for State-of-the-Art (SOTA) models. However, ChatGPT was not evaluated on the IE dataset due to potential data leakage concerns. In contrast, Llama-3.3-70B-Instruct was run locally, allowing for GPU resource estimation and comprehensive metric evaluation. As presented in Table 3, the highest prompt engineering performance was achieved by ChatGPT-o1 mini. Meanwhile, in the IE dataset, which serves as our industrial database, an accuracy score of 0.800 was achieved by Llama-3.3-70B-Instruct. Regarding SLMs, we found the results particularly intriguing, especially in the context of prompt engineering. Given the relatively small size of these models, we did not expect them to achieve high performance. The final results for the 1B and 3B models aligned with our expectations, performing roughly at the level of random guessing. Interestingly, both the 3B and even the 1B models demonstrated a strong preference for few-shot prompting. This approach led to an improvement of over 10% compared to the base prompt on the EUR and LDD datasets, highlighting the importance of few-shot learning in the application of SLMs, as also emphasized in Brown et al. (2020a). Furthermore, we observed that both COD and COT provided limited improvements. In fact, on the LDD dataset, COD performed worse than COT and was nearly on par with the base prompt. Therefore, the use of COD and COT is not recommended as a solution for classification tasks in SLMs.
In the context of SFT, we observed that SPT outperformed prefix tuning by a significant margin, although it also required substantially more training time. Prefix tuning introduces a trainable part at every layer within the model, whereas SPT only incorporates a soft prompt at the input level. It is possible that SPT better preserves the original language understanding of the model, as it does not alter the overall architecture. In contrast, prefix-tuning’s modifications to the attention structure may disrupt the model’s inherent linguistic comprehension. Additionally, supervised fine-tuning, which adds a classification head to the end of the model, demonstrated the highest overall performance. Notably, ModernBERT achieved a performance of approximately 0.810 of accuracy on the LDD dataset while requiring less training time and GPU memory, making it a promising candidate for academic English text classification. Limited exposure to French, other multilingual languages, and domain-specific corpora during training Warner et al. (2024) led to weaker performance on the IE dataset (primarily in French) and EUR (a domain-specific corpus).
@@ 4.2.3 Deeper Header?
In our primary experimental setting, we adhere to the definition of a “Header” as implemented in the Transformers library, referring to a single linear layer serving as the classification head. To further explore potential improvements using different levels of header, we experimented with replacing the standard single-layer header with a multi-layer linear architecture incorporating ReLU activations. Specifically, we constructed classification heads with 2 to 5 linear layers (hidden dimension = 256) and fine-tuned Llama-3.2-1B-Instruct model accordingly. As shown in Table 5, the results indicate that increasing the depth of the classification head yields only marginal gains, with performance plateauing beyond three layers. These findings suggest that deeper header architectures offer limited benefit in enhancing the classification accuracy or F1 score in this context.
@@ 4.4 Research Questions
For RQ1, “Can SLMs perform classification without any task-specific training?”, we found that text classification using SLMs faces several key challenges. Smaller models tend to exhibit limited logical reasoning capabilities and are more susceptible to generating hallucinations while encountering long text. Moreover, the performance ceiling is strongly influenced by the amount of available training data, while the intrinsic properties of the SLMs themselves also play a critical role in shaping classification outcomes.
@@
For RQ3, “How can the trade-off between computational efficiency and classification performance be optimized, and how can SLMs be more effectively deployed in practice?”, we found that although training the model consumes significant GPU resources, the SLMs are essentially unusable in their current form due to the lack of inference capability. We also tested Llama-3.3-70B-Instruct, which, although capable of achieving 80% accuracy in IE, still produces uncertain output. Therefore, FT transformers remains the only viable solution on SLMs which is portable and light weight. Finally, the limited capacity of SLMs creates a bottleneck on performance and the amount of labeled data also remains a key limitation. For real application, it is crucial to focus not only on data quality but also on the model’s inherent characteristics, such as multilingual comprehension. If resources are relatively abundant, opting for decoder-only models such as the Llama series would be a better choice, which has a good support on both languages and different domain knowledge.
@@ 5 Conclusion
In this study, we present a comprehensive evaluation of lightweight models on text classification. We systematically investigate nearly all major approaches, including prompt engineering and supervised fine-tuning. Our experimental setup spans three benchmark datasets, including a real-world industrial scenario involving email history classification.